OpenStudy (anonymous):

how do i convert 1 3/4

OpenStudy (anonymous):

convert to a percent or decimal?

OpenStudy (paki):



OpenStudy (anonymous):

decimal is .75

OpenStudy (anonymous):

\[ \huge 1\tfrac{3}{4}=\frac{4\times 1+3}{4}=\frac{7}{4}\]